OpenStudy (anonymous):

I need a formula please! Here is the question: you invest $2200 at an interest rate that is compounded annually. After 8 years, the principal amount has increased to $2896.98. What percent is the interest rate?

OpenStudy (anonymous):

i can calculate compound interest WITH a rate, but not sure how to do it backwards! Help!

OpenStudy (amistre64):

A = P (1+r)^t A/P = (1+r)^t t.rt(A/P) = 1+r t.rt(A/P) -1 = r maybe?

OpenStudy (amistre64):

8rt(2896.98/2200) - 1 might do it

OpenStudy (amistre64):

3.5% if soo

OpenStudy (amistre64):

if 2200 (1.035)^8 = the larger amount were good

OpenStudy (anonymous):

Again a bit of confusion - one more question!!

OpenStudy (amistre64):

its good

OpenStudy (anonymous):

OK - that's what I was confused about...the 8 is the exponent to the 1.035, correct?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

since its compounded 1 a year, there is no reason to do anything fancier

OpenStudy (anonymous):

Thanks! That makes way more sense!

OpenStudy (amistre64):

A = P (1+r)^t A/P = (1+r)^t ln(A/P) = t ln(1+r) yeah, the only way i see to get to the r is to take the "t"root of it all